On January 1, 2021, Gundy Enterprises purchases an office building for $283,000, paying $53,000 down and borrowing the remaining $230,000, signing a 8%, 10-year mortgage. Installment payments of $2,790.53 are due at the end of each month, with the first payment due on January 31, 2021.

Record the purchase of the building on January 1, 2021. (If no entry is required for a particular transaction/event, select "No Journal Entry Required" in the first account field.)

Complete the first three rows of an amortization schedule. (Do not round intermediate calculations. Round your final answers to 2 decimal places.)

Record the first monthly mortgage payment.

Total payments over the 10 years are $334,864 ($2,790.53 × 120 monthly payments). How much of this is interest expense and how much is actual payment of the loan?
Answer #1

1. 01/01/2021 Debit Office Account 283000

Credit Cash 53000

Credit Note Payable 230000

2. Amortization payment for every month given by formula = E = P * r * (((1+r)^n) / (((1+r)^n) -1))

E = 230000*(0.08/12)*((1+ 0.08/12)^120 / (1+ 0.08/12)^120 -1)) = 2790.53

For first month:

Interest component = P*r/12 = 230000*0.08/12 = 1533.33

Principal component = E-Principal component = 2790.53 - 1533.33 =1257.20

for Second month : Principal get subtracted by Principal component paid in first month=>

=> P = 230000 - 1257.20 = 228742.8

In Co = 228742.8*(0.08/12) = 1524.95

Pr Co = 2790.53 - 1524.95 = 1265.58

for Third month : second month Principal get subtracted by Principal component paid in second month

=>P = 228742.8-1265.58 = 227477.22

In Co = 227477.22*(0.08/12) = 1516.52

Pr Co = 2790.53 - 1516.52 = 1274.01

Date Interest Principal Balance
Jan, 2021 $1,533 $1,257 $228,743
Feb, 2021 $1,525 $1,266 $227,477
Mar, 2021 $1,517 $1,274 $226,203

Out of total payment of $334,864:

Principal paid = $230,000

Interest paid = $334864 - $230000 = $104,864